OpenStudy (anonymous):

Use power series to solve the differential equation \[y''+xy'+y=0\] \[y=\sum_{n=0}^{\infty} a_n x^n\] \[y'=\sum_{n=1}^{\infty}a_nnx^{n-1}\] \[y''=\sum_{n=2}^{\infty} n(n-1)a_nx^{-2}\] \[xy'=\sum_{n=1}^{\infty}a_nnx^n\] \[\sum_{n=2}^{\infty} n(n-1)a_nx^{-2}+\sum_{n=1}^{\infty}a_nnx^n+\sum_{n=0}^{\infty} a_n x^n=0\]

OpenStudy (turingtest):

yep, now what's first?

OpenStudy (anonymous):

changing the sums so that they all begin with n=2 changing the last sum from n=0 to n=1 would be \[a_1+\sum_{n=1}^\infty a_nx^n\]

OpenStudy (turingtest):

no we don't strip out the terms yet, we just do an index shift

OpenStudy (turingtest):

the first order of business is to get all the x's to be to the nth power through using index shifts

OpenStudy (turingtest):

the only series that has x to some power other than n is the first one how can we change the index of the first series to make the exponent on x be n?

OpenStudy (anonymous):

you mean on \[y''=\sum_{n=2}^{\infty} n(n-1)a_nx^{n-2}\]

OpenStudy (turingtest):

yes, we need to start from a different n to get x^n through an index shift

OpenStudy (anonymous):

so I'm shifting the starting point up by 2 so I have to subtract 2 from {n=2-2} \[-\sum_{n=0}^\infty n(n-1)a_nx^n\] negative to make up for the negative?

OpenStudy (anonymous):

nooooo

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

\[\sum_{n=0}^\infty n(n+1)a_nx^n\]

OpenStudy (turingtest):

nice :)

OpenStudy (anonymous):

still not right

OpenStudy (anonymous):

because that would give us zero

OpenStudy (turingtest):

no but closer

OpenStudy (turingtest):

you didn't add 2 to every n

OpenStudy (anonymous):

\[\sum_{n=0}^\infty (n+2)(n+1)a_nx^n\]

OpenStudy (turingtest):

and I do mean \(every\) n (subscripts included)

OpenStudy (turingtest):

again better, but...^

OpenStudy (anonymous):

\[\sum_{n=0}^\infty (n+2)(n+1)a_{n+2}x^n \]

OpenStudy (anonymous):

can I add them?

OpenStudy (turingtest):

now that, I do believe, is correct :)

OpenStudy (anonymous):

or multiply

OpenStudy (anonymous):

or just leave as is?

OpenStudy (turingtest):

add what? oh that? I wouldn't yet

OpenStudy (anonymous):

ok

OpenStudy (turingtest):

yeah leave as is asd now rewrite what you have

OpenStudy (turingtest):

and*

OpenStudy (turingtest):

the whole expression please I'm lazy

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty} (n+2)(n-1)a_{n+2}x^{n}+\sum_{n=1}^{\infty}a_nnx^n+\sum_{n=0}^{\infty} a_n x^n=0\]

OpenStudy (turingtest):

\[\sum_{n=0}^{\infty} (n+2)(n+1)a_{n+2}x^{n}+\sum_{n=1}^{\infty}a_nnx^n+\sum_{n=0}^{\infty} a_n x^n=0\]okay next order of business is what you tried to do earlier; get all indices to start at the same n value the only problem child here is which term?

OpenStudy (anonymous):

the middle

OpenStudy (turingtest):

so what do we do to fix it?

OpenStudy (turingtest):

you have to be a bit clever here...

OpenStudy (anonymous):

\[ xa_n +\sum_{n=0}^\infty a_n(n+1)x^n \]

OpenStudy (turingtest):

did you try to do an index shift?

OpenStudy (turingtest):

it looks like you combined the idea of the index shift with stripping out the first term

OpenStudy (anonymous):

oh my...let's try this again

OpenStudy (turingtest):

it is a bit tricky so let me give you a hint...

OpenStudy (turingtest):

you cannot do another index shift. If you do you will change the exponent on x, which means we are screwed again, so we need to do something else to get this to start from n=0

OpenStudy (turingtest):

we can't really strip out a term because we don't have this series defined for n<1, but ask yourself, if it was defined at n=0, what would the first term be?

OpenStudy (anonymous):

just have the addition of an a_n up front like in the last example?

OpenStudy (turingtest):

no, answer my question above and you will see why...

OpenStudy (anonymous):

a 1

OpenStudy (anonymous):

times a_n

OpenStudy (turingtest):

no, be more careful

OpenStudy (turingtest):

what would be the first term of\[\sum_{n=0}^\infty a_n nx^n\]

OpenStudy (anonymous):

zero

OpenStudy (turingtest):

right, so...

OpenStudy (turingtest):

\[0+\sum_{n=1}^\infty a_n nx^n=\sum_{n=0}^\infty a_n nx^n\]so we actually didn't need to change anything; the first term we "stripped out" was a zero anyway! It makes no difference.

OpenStudy (turingtest):

so we can go ahead and just change the starting point of that series without further manipulation

OpenStudy (turingtest):

always watch for that, it comes in handy

OpenStudy (anonymous):

oh ok uhm.....I'll revisit the last example later, because the I'm not convinced that we needed to do a index shift (or was it stripping of a term)

OpenStudy (turingtest):

first to get the x^n on all terms we did an index shift (where we add or subtract from each n in the summand) next to get all the starting points for each series the same, we strip out terms/or do the trick we just did they are different tricks, and the order in which you do them is imortant

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty} (n+2)(n+1)a_{n+2}x^{n}+\sum_{n=0}^{\infty}a_nnx^n+\sum_{n=0}^{\infty} a_n x^n=0\]

OpenStudy (turingtest):

it is important for you to try to understand why the index shift requires you to add/subtract each n in the summand while stripping out a term does not

OpenStudy (turingtest):

good, so now make it all into one summation...

OpenStudy (anonymous):

\[\sum_{n=0}^\infty x^n[(n+2)(n+1)a_{n+2}+a_nn+a_n]=0\]

OpenStudy (turingtest):

nice, so what will you set to zero to find the recurrence relation?

OpenStudy (turingtest):

(I am going to type minimally now as I am eating dinner at the same time :)

OpenStudy (anonymous):

food comes first in my book... :P Ok let's see here...

OpenStudy (turingtest):

I can eat and respond with "yes" or "no" pretty well, no worries :)

OpenStudy (anonymous):

\[(n+2)(n+1)a_{n+2}+a_nn+a_n=0\implies a_{n+2}=-\frac{a_nn-a_n}{(n+2)(n+1)}\]

OpenStudy (anonymous):

si?

OpenStudy (turingtest):

si, pero si quieras podrias factor la \(-a_n\)

OpenStudy (anonymous):

ok \[a_{n+2}=-a_n\frac{n-1}{(n+2)(n+1)}\]

OpenStudy (turingtest):

and now comes the tedious part that you can do while I eat plug in n=0,1,2,3 and check what results from that recurrence relation until you notice a pattern like last time so just cranks it out; plug in n=0,n=1,n=2 and if a_2 can be written in terms of a_0 or something like last time, do it. That is how we find the pattern.

OpenStudy (turingtest):

go ahead and list out the first 5 or so and let's see if we see a relation

OpenStudy (anonymous):

\[n=1:a_3=0\] \[n=2:a_4=-a_2\frac{1}{12}\] \[n=3:a_5=-a_3\frac{2}{20}\] \[n=4:a_6=-a_4\frac{3}{30}\] \[n=5:a_7=-a_5\frac{4}{42}\]

OpenStudy (turingtest):

what about n=0 ? we need to start from n=0 since our series starts from n=0 this time (last time it started from n=1)

OpenStudy (anonymous):

\[n=0:a_2=-a_0\frac{-1}{2}\]

OpenStudy (anonymous):

\[n=0:a_2=a_0\frac{1}{2}\]

OpenStudy (anonymous):

whatcha eating?

OpenStudy (turingtest):

torta de pollo=chicken sandwich on a french sort of bread

OpenStudy (turingtest):

con chipotle

OpenStudy (anonymous):

nice

OpenStudy (anonymous):

oh yeah...the pattern

OpenStudy (turingtest):

so you did well in your subs I think, but now you must try to always sub all the way back as far as you can..

OpenStudy (turingtest):

yeah, the pattern write each on as some finite number of unknowns, in this case everything in terms of \(a_0\) and \(a_3\)

OpenStudy (turingtest):

each one*

OpenStudy (anonymous):

in the previous example you seemed to have simply switched the \[a_{odd}\] to zero and the \[a_{even}\] to \[a_0\]

OpenStudy (anonymous):

take you're time...i'm gonna watch a few more mins of my episode

OpenStudy (turingtest):

it's not different, it's the same as last time odd subscript gives zero even has a pattern reducible to a_0

OpenStudy (turingtest):

okay I got a fairly simple answer, let me see if I can confirm it...

OpenStudy (turingtest):

stupid wolfram is of no help :/

OpenStudy (anonymous):

\[n=0:a_2=a_0\frac{1}{2}=\frac{a_0}{2}\] \[n=1:a_3=0\] \[n=2:a_4=-a_2\frac{1}{12}=\frac{a_0}{2(3!)}\] \[n=3:a_5=0\] \[n=4:a_6=-a_4\frac{3}{30}=\frac{a_0}{5(3!)}\] \[n=5:a_7=0\]

OpenStudy (anonymous):

how did I do?

OpenStudy (turingtest):

almost, I actually realized that I made a mistake because I think I see a mistake from you so you did as well as me I'd say :D

OpenStudy (turingtest):

let me try on paper again...

OpenStudy (turingtest):

here I think I can demonstrate the benefit of not multiplying out terms and writing things over-explicitly to find patterns...

OpenStudy (anonymous):

ok

OpenStudy (turingtest):

\[n=0:a_2=\frac{a_0}{2\cdot 1}\]\[n=2:a_4\frac{a_2}{4\cdot 3}=\frac{a_0}{4\cdot3\cdot2\cdot 1}\]now reevaluate the pattern

OpenStudy (anonymous):

\[n=0:a_2=\frac{a_0}{2\cdot 1}=\frac{a_0}{2!}\] \[n=2:a_4\frac{a_2}{4\cdot 3}=\frac{a_0}{4\cdot3\cdot2\cdot 1}=\frac{a_0}{4!}\]

OpenStudy (turingtest):

so try to do the same thing as last time name all even numbers and make a new series with index k

OpenStudy (turingtest):

then represent the coefficients and plug that in for \(a_n\)

OpenStudy (anonymous):

n=2k \[a_n=a_{2k}=\frac{a_0}{k!}\]

OpenStudy (turingtest):

is it over k! ?